OpenStudy (anonymous):

What exponential function is the best fit for the data in the table? x f(x) 1 -4 3 -1 4 3 f(x) = 4(3)x - 1 + 4 f(x) = 4(3)x - 1 - 4 f(x) = one fourth(3)x - 1 + 4 f(x) = one fourth(3)x - 1 - 4

OpenStudy (anonymous):

f(x) = 4(3)x - 1 + 4 f(x) = 4(3)x - 1 - 4 f(x) = one fourth(3)x - 1 + 4 f(x) = one fourth(3)x - 1 - 4 Answer choices ^^

OpenStudy (ilovecake):

I would say the last one cause that one is the only answer that works. SO it is f(x) = one fourth(3)x - 1 - 4

OpenStudy (anonymous):

Thank you so much can you help me with one more?

OpenStudy (ilovecake):

ok

OpenStudy (ilovecake):

btw sorry I answered so late

OpenStudy (anonymous):

No problem at all !!!! The next one is ... Which graph models the function f(x) = −4(2)x?

OpenStudy (ilovecake):

Do you have any answer choices?

OpenStudy (anonymous):

its -4(2)^x

OpenStudy (ilovecake):

Oh.....

OpenStudy (ilovecake):

ok hold on

OpenStudy (ilovecake):

click on the link

OpenStudy (anonymous):

My computer is not working well and that's why I'm taking long to respond

OpenStudy (ilovecake):

That gives the answer

OpenStudy (ilovecake):

Okay. But, did you see what I sent you?

OpenStudy (ilovecake):

Sorry about that

OpenStudy (anonymous):

That isn't an answer choice.

OpenStudy (anonymous):

Yeah. My computer's not letting me post the answer choices.

OpenStudy (ilovecake):

I know. But, it helps you to find the answer

OpenStudy (ilovecake):

I will fan you. Then you could sent it through the messages

OpenStudy (anonymous):

It's okay. I got it. Thank you for helping.

OpenStudy (ilovecake):

Oh ok.

OpenStudy (ilovecake):

Good luck